OpenStudy (anonymous):

HELP! For f(x)=xsqrt|x} find the maximum domain, maximum continuity range, and the range of the derivative.

OpenStudy (anonymous):

\(f(x)=x\sqrt{x}\)

OpenStudy (anonymous):

is that it? or something else?

OpenStudy (anonymous):

\[f(x)=x \sqrt{|x|}\]

OpenStudy (anonymous):

oh that is different first the domain, since \(|x|\geq 0\) always, the domain is all real numbers

OpenStudy (anonymous):

then there are two choices, \(x\geq 0\) in which case this is \(x\sqrt{x}\) or \(x^{\frac{3}{2}}\) otherwise if \(x<0\) this would be \(x\sqrt{-x}\)

OpenStudy (anonymous):

you are talking about the domain or continuity? I thought you said domain was for all real numbers so why must it be split up to positive and negative?